# ColorwAI: Generative Colorways of Textiles through GAN and Diffusion Disentanglement

### v1

@@ Appendix A Diffusion Denoising Probabilistic Models
(5) $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$
@@
(6) $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$
@@
DiffusionCLIP (Kim et al., 2022) integrates diffusion models with Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) to guide the diffusion process, enabling the generation of images that are semantically aligned with provided textual descriptions. Let $\mathbf{y}$ be the textual description and $\mathbf{e}_{y}$ be the corresponding text embedding obtained from CLIP. The diffusion model’s reverse process is then conditioned on this text embedding $\mathbf{e}_{y}$:
(7) $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{e}_{y})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t,\mathbf{e}_{y}),\Sigma_{\theta}(\mathbf{x}_{t},t,\mathbf{e}_{y}))$
Asyrp (Kwon et al., 2022) propose to modify the generative reverse process to use the deepest feature maps, the h-space, of a pre-trained diffusion model as semantic latent space for image manipulation. DDIM (Song et al., 2020) redefines Eq. 1 as
(8) $q_{\sigma}(x_{t-1}|x_{t},x_{0})=\mathcal{N}\left(\sqrt{\alpha_{t-1}}x_{0}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\frac{x_{t}-\sqrt{\alpha_{t}}x_{0}}{\sqrt{1-\alpha_{t}}},\sigma_{t}^{2}\mathbf{I}\right)$
@@
(9) $x_{t-1}=\sqrt{\alpha_{t-1}}P_{t}(\epsilon_{\theta}(x_{t}|\Delta h_{t}))+D_{t}+\sigma_{t}z_{t},$
@@
(10) $L_{\text{direction}}(x_{\text{edit}},y_{\text{target}};x_{\text{source}},y_{\text{source}})=1-\frac{\Delta I\cdot\Delta T}{\|\Delta I\|\|\Delta T\|},$
@@
Boundary Diffusion (Zhu et al., 2023) enables precise attribute control by directly segmenting the latent space into distinct regions without adding any extra parameters to the base diffusion model. This is in contrast to Asyrp, which requires training auxiliary editing neural networks, and DiffusionCLIP, which fine-tunes the pre-trained diffusion model. Boundary Diffusion introduces a modified reverse process of eq. 9 to include boundaries in the latent space:
(11) $x_{t-1}=\sqrt{\alpha_{t-1}}P_{t}(\epsilon_{\theta}(x_{t}|B_{t}(\Delta h_{t})))+D_{t}+\sigma_{t}z_{t}$
@@
BoundaryDiffusion finds the disentangled direction of a feature by obtaining the semantic boundary of that feature using SVM. The latent direction is the orthonormal direction exiting the semantic boundary. The method to find the disentangled direction is equivalent to the InterfaceGAN proposed for GANs (Shen et al., 2020). The applicability of this method on the latent space of Diffusion indicates that, within the BoundaryDiffusion framework, also the other GAN disentanglement methods, StyleSpace and ShapleyVec, can be applied.
Appendix B Color annotation
B.0.1. Color Annotation
As we wish to disentangle coarse color directions, that would allow harmonious modifications based on a predominant guiding color, we annotate the predominant color of each image. Given a latent code $\mathbf{z}$ that generates an image $\mathbf{x}$, we annotate the main color $\mathbf{s}$ of $\mathbf{x}$. Digitally, color can be represented in different encodings, notably, RGB, HSV, CIELab, XYZ, each enforcing an inherent concept of similarity and gamuts. Particularly, the encoding of CIELab is most aligned to human vision, while HSV (hue, saturation, value), is the most understandable and intuitive color representation (Mojsilovic et al., 2000).
@@ B.0.1. Color Annotation
Several methods have been developed to extract palettes from an image, often in the domains of image compression or retrieval. The most simple method relies on k-means to identify a set of $K$ clusters from the $RGB$ or $Lab$ colors of an image and take the cluster centroids to be the palette. While this method works well, it fails to extract contrasting colors, which stand out the most to the human eye. Modifications have been proposed, including adaptive clustering methods (Pappas, 1989; Bhatia et al., 2004), and codebook-based quantization methods (Mojsilovic et al., 2000). We test the different methods on our data and find empirically that an adaptation of (Bhatia et al., 2004)’s method using $CIELab$ encoding results in the most pleasing and varied color palette.
@@
Figure 5. Predominant color annotation. From left to right: palette extraction, codebook identification, color quantization. The final annotated color is the lower box of each rightmost image.
[…]
Predominant color annotation. From left to right: palette extraction, codebook identification, color quantization. The final annotated color is the lower box of each rightmost image.

### v2

@@ Appendix 0.A Diffusion Denoising Models
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$ (5)
@@
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$ (6)
@@
DiffusionCLIP [23] integrates diffusion models with Contrastive Language-Image Pre-training (CLIP) [36] to guide the diffusion process, enabling the generation of images that are semantically aligned with provided textual descriptions. Let $\mathbf{y}$ be the textual description and $\mathbf{e}_{y}$ be the corresponding text embedding obtained from CLIP. The diffusion model’s reverse process is then conditioned on this text embedding $\mathbf{e}_{y}$:
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{e}_{y})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t,\mathbf{e}_{y}),\Sigma_{\theta}(\mathbf{x}_{t},t,\mathbf{e}_{y}))$ (7)
Asyrp [26] propose to modify the generative reverse process to use the deepest feature maps, the h-space, of a pre-trained diffusion model as semantic latent space for image manipulation. DDIM [42] redefines Eq. 5 as
$q_{\sigma}(x_{t-1}|x_{t},x_{0})=\mathcal{N}\left(\sqrt{\alpha_{t-1}}x_{0}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\frac{x_{t}-\sqrt{\alpha_{t}}x_{0}}{\sqrt{1-\alpha_{t}}},\sigma_{t}^{2}\mathbf{I}\right)$ (8)
@@
$x_{t-1}=\sqrt{\alpha_{t-1}}P_{t}(\epsilon_{\theta}(x_{t}|\Delta h_{t}))+D_{t}+\sigma_{t}z_{t},$ (9)
@@
$L_{\text{direction}}(x_{\text{edit}},y_{\text{target}};x_{\text{source}},y_{\text{source}})=1-\frac{\Delta I\cdot\Delta T}{\|\Delta I\|\|\Delta T\|},$ (10)
@@
Boundary Diffusion [51] enables precise attribute control by directly segmenting the latent space into distinct regions without adding any extra parameters to the base diffusion model. This is in contrast to Asyrp, which requires training auxiliary editing neural networks, and DiffusionCLIP, which fine-tunes the pre-trained diffusion model. Boundary Diffusion introduces a modified reverse process of eq. 9 to include boundaries in the latent space:
$x_{t-1}=\sqrt{\alpha_{t-1}}P_{t}(\epsilon_{\theta}(x_{t}|B_{t}(\Delta h_{t})))+D_{t}+\sigma_{t}z_{t}$ (11)
@@
BoundaryDiffusion finds the disentangled direction of a feature by obtaining the semantic boundary of that feature using SVM. The latent direction is the orthonormal direction exiting the semantic boundary. The method to find the disentangled direction is equivalent to the InterfaceGAN proposed for GANs [41]. The applicability of this method on the latent space of Diffusion indicates that, within the BoundaryDiffusion framework, also the other GAN disentanglement methods, StyleSpace and ShapleyVec, can be applied.
Appendix 0.B Color annotation
0.B.0.1 Color Annotation
As we wish to disentangle coarse color directions, that would allow harmonious modifications based on a predominant guiding color, we annotate the predominant color of each image. Given a latent code $\mathbf{z}$ that generates an image $\mathbf{x}$, we annotate the main color $\mathbf{s}$ of $\mathbf{x}$. Digitally, color can be represented in different encodings, notably, RGB, HSV, CIELab, XYZ, each enforcing an inherent concept of similarity and gamuts. Particularly, the encoding of CIELab is most aligned to human vision, while HSV (hue, saturation, value), is the most understandable and intuitive color representation [32].
@@ 0.B.0.1 Color Annotation
Several methods have been developed to extract palettes from an image, often in the domains of image compression or retrieval. The most simple method relies on k-means to identify a set of $K$ clusters from the $RGB$ or $Lab$ colors of an image and take the cluster centroids to be the palette. While this method works well, it fails to extract contrasting colors, which stand out the most to the human eye. Modifications have been proposed, including adaptive clustering methods [33, 4], and codebook-based quantization methods [32]. We test the different methods on our data and find empirically that an adaptation of [4]’s method using $CIELab$ encoding results in the most pleasing and varied color palette.
@@
Figure 5: Predominant color annotation. From left to right: palette extraction, codebook identification, color quantization. The final annotated color is the lower box of each rightmost image.
Appendix 0.C Questionnaire
This appendix documents the questions and visuals of the expert questionnaire.
  1. 1.
    Introduction
    • Name
    • Email
    • Job position and company/institution
    • On a scale from 1 to 5, how do you evaluate your level of experience with design and textiles? (1 not expert, 5 expert)
  2. 2.
    Creativity
    • Novelty. On a scale of 1 to 5, how innovative/original do you find the color combinations generated compared to existing designs? (1 not new at all, 5 extremely new)
    • If you think one or more combinations are innovative, indicate which ones and explain why
    • Surprise. On a scale of 1 to 5, how surprising do you rate the color and pattern combinations generated? (1 not at all surprising, 5 extremely surprising)
    • Based on the examples presented, how would you define the type of surprise they elicit?
    • Usefulness. On a scale of 1 to 5, how applicable are the examples of "generative colorway" to possible commercial textile collections? (1 not useful at all, 5 extremely useful)
    • Why?
  3. 3.
    Usability
    • Ease of interaction. On a scale of 1 to 5, how do you find the choice of color modification based only on a predominant color and not on the entire palette? (1 very difficult, 5 very easy)
    • What advantages and disadvantages do you identify?
    • Control and flexibility. On a scale of 1 to 5, do you feel that you can have sufficient control over the colors using this simplified model? (1 no control, 5 complete control)
    • What improvements could be introduced to increase control and/or flexibility?
    • Color choice. On a scale of 1 to 5, how comprehensive do you find the spectrum of colors available to you? (1 not exhaustive at all, 5 comprehensive)
    • Assuming, for example, that you need to rework a historical textile, what color is missing? How many color choices would you like to have (there are currently 19)?
  4. 4.
    Aesthetics
    • Visual attractiveness. On a scale of 1 to 5, how aesthetically pleasing do you find the generated color combinations? (1 not pleasing at all, 5 extremely pleasing)
    • What specific elements of color combinations contribute to greater appeal?
    • Shape modifications. On a scale of 1 to 5, how do slight pattern adjustments contribute to the overall aesthetic quality of the fabric? (1 detrimental, 5 significantly improving)
    • Are there specific adjustments that you have found particularly effective or not?
    • Color combination. On a scale of 1 to 5, how harmonious are the colors used in the textile? (1 not harmonious at all, 5 extremely harmonious)
    • What color combinations do you like best and why?
  5. 5.
    Consistency with textile design practices and traditions
    • Consistency. On a scale of 1 to 5, do color combinations and patterns maintain consistency with established design principles? (1 not consistent at all, 5 highly consistent)
    • In what examples, do you find that more or less of these principles have been adhered to? Which ones?
    • Historical Pertinence. On a scale of 1 to 5, how consistent are the color palette combination and pattern with the historical period from which the pattern came? (1 not accurate at all, 5 highly accurate)
    • Among the examples, are there any styles or historical periods with which the AI color combinations would be assonant or not?
  6. 6.
    Coherence between pattern and color
    • Necessity of the pattern change. On a scale of 1 to 5, how necessary do you think the pattern adjustments made by the AI to accommodate the new color scheme are? (1 not necessary at all, 5 extremely necessary)
    • Explain why you think these adjustments are necessary or unnecessary.
    • Improvement in consistency. On a scale of 1 to 5, how do these minimal changes in shape impact pattern-color consistency? (1 significant worsening, 5 significant improvement)
    • Among the examples, are there instances where consistency was particularly strengthened or not?
  7. 7.
    General feedback and suggestions
    • General impression. What is your overall impression of ""generative colorway""? Are there any aspects that you particularly liked or disliked?
    • Suggestions for improvement. Do you have suggestions for improving AI performance in ""generative colorway""? Are there additional features or controls that you would like to see integrated into the system?
[Uncaptioned image][Uncaptioned image]
Refer to captionRefer to captionRefer to caption
Figure 6: Examples of "generative colorways" to be considered in each section. On the left is the image of the original pattern; on the right are three versions of "generative colorway" in the guide color indicated by the arrow.